### v5

@@ IV.2 Critical time scale
In the ergodic case, the setting of $\tau^{\left(q\right)}$ in the SRG (as we have mentioned in step (2)) is finished in $1$-st iteration and required to make the specific heat, an indicator of the transitions of diffusion processes, maximize (see similar ideas in Ref. [10]). In Appendix C, we explain why it is valid to define renormalization flows based on the specific heat. Specifically, we consider the $q$-order spectral entropy in the $1$-st iteration of the SRG (see Ref. [42] for the definition of spectral entropy)
@@
where $\big{\langle}\lambda^{\left(q\right)}_{1}\big{\rangle}_{\rho}=\operatorname{tr}\left(\mathbf{L}^{\left(q\right)}_{1}\rho^{\left(q\right)}_{1}\right)$. The $q$-order specific heat is defined according to the first derivative of the $q$-order spectral entropy
@@
Please see Appendix D for detailed derivations. The value of $\tau^{\left(q\right)}$ is selected to ensure $\frac{\mathsf{d}}{\mathsf{d}\tau^{\left(q\right)}}X^{\left(q\right)}_{1}\left(\tau^{\left(q\right)}\right)=0$, which indicates an infinite deceleration condition of diffusion processes. See Fig. 3(c) for illustrations. Apart from the infinite deceleration condition, determining the time scale value $\widehat{\tau}^{\left(q\right)}$ used in the step (2) of the SRG also requires us to consider the high-order properties of diffusion. For convenience, we denote $\widetilde{\tau}^{\left(q\right)}$ as the value of $\tau^{\left(q\right)}$ under the infinite deceleration condition (i.e., there is $\frac{\mathsf{d}}{\mathsf{d}\tau^{\left(q\right)}}X^{\left(q\right)}_{1}\left(\tau^{\left(q\right)}\right)\big{|}_{\widetilde{\tau}^{\left(q\right)}}=0$). The optimal time scale for the SRG is set as
$\widehat{\tau}^{\left(q\right)}=\frac{2}{q+1}\widetilde{\tau}^{\left(q\right)},$ (24)
@@
which is the average subdivided time scale for each $1$-simplex in a $n$-simplex (i.e., the average number of adjacent $1$-simplexes between arbitrary units $i$ and $j$ is $\frac{n+1}{2}$). Note that $\widehat{\tau}^{\left(q\right)}=\widetilde{\tau}^{\left(q\right)}$ holds for pairwise interactions.
@@
The above procedure can be directly generalized to the non-ergodic case, where we treat each cluster $C_{i}$ as a network and derive a $\widetilde{\tau}^{\left(q\right)}$ for it if it has not been assigned with a time scale yet. The time scale of a cluster does not change after it is assigned. See Fig. 3(c) for illustrations.
After determining the time scale, the SRG proposed in Sec. IV.1 can be applied to renormalize interacting systems in a non-parametric manner. In Appendix E, we use the SRG to process a system whose pairwise interactions exhibit small-world properties [80] and illustrate how the renormalization flow changes across different combinations of $\left(p,q\right)$. This result qualitatively suggests the effects of higher-order interactions on lower-order interactions during renormalization, whose quantification is explored in our subsequent analysis. In Fig. 4, we present more instances of the SRG applied on synthetic interacting systems whose pairwise interactions follow the random tree (RT), the square lattice (SL), the triangular lattice (TL), and the Watts-Strogatz networks (WS). Please see Appendix A for the definitions of these random networks.
IV.3 Scaling relation and high-order effects
@@
where $\big{\langle}\lambda^{\left(q\right)}_{1}\big{\rangle}_{\widetilde{\rho}}=\operatorname{tr}\left(\mathbf{L}^{\left(q\right)}_{1}\widetilde{\rho}^{\left(q\right)}\right)$ and constant $\mu$ is obtained when we solve Eq. (23) as a differential equation. If the Laplacian eigenvalue spectrum (i.e., the density of the eigenvalues of $\mathbf{L}^{\left(q\right)}_{1}$) follows a general power-law form $\operatorname{Prob}\left(\lambda^{\left(q\right)}_{1}\right)\sim\left(\lambda^{\left(q\right)}_{1}\right)^{\beta^{\left(q\right)}}$, we can obtain a scaling relation
@@
whose derivations are presented in Appendix F. Parameter $\nu$ is a specific constant. For a system with scale-invariance property on the $q$-order, the power-law form of the Laplacian eigenvalue spectrum holds and should be invariant under the transformation of the SRG. Therefore, we expect to see an approximately constant $\widetilde{X}^{\left(q\right)}$ in a certain interval of time scale (i.e., the infinite deceleration condition), accompanied by a fixed $\beta^{\left(q\right)}$ during renormalization. For a system without scale-invariance, the scaling relation vanishes.
Figure 5: Scaling relation of the simplex renormalization group. (a-e) We verify the scaling relation on the systems whose pairwise interactions follow a Barabási-Albert network (BA, $c=10$), a Watts-Strogatz network (WS, each unit initially has $10$ neighbors and edges are rewired according to a probability of $0.1$), and an Erdos-Renyi network (ER, two units share an edge with a probability of $0.2$), the triangular lattice (TL, with the interactions of the first two orders), and the random tree (RT, with only pairwise interactions), respectively. Scatters denote the mean values of $\left(\widetilde{X}^{\left(q\right)},\widetilde{\tau}^{\left(q\right)}\Big{\langle}\frac{\beta^{\left(q\right)}\left(i\right)+1}{\widetilde{\tau}^{\left(q\right)}\left(i\right)}\Big{\rangle}_{i}+\nu\right)$ averaged across all replicas on each order and error bars denote standard deviations. See the definitions of random networks in Appendix A. (f) The standard deviations compared across different systems. (g-i) The high-order effects measured on the replicas in (a-c), where each color area denotes the distribution of high-order effects in the $k$-iteration of the SRG ($k\in\{1,2,3\}$).
To generalize our analysis to the non-ergodic case, we respectively derive $\widetilde{\tau}^{\left(q\right)}\left(i\right)$ and $\beta^{\left(q\right)}\left(i\right)$ for each cluster $C_{i}$ of $\mathbf{G}^{\left(q\right)}_{1}$ following Eqs. (25-26). These definitions enable us to obtain
@@
where $\langle\cdot\rangle_{i}$ denotes weighted averaging across all clusters of $\mathbf{G}^{\left(q\right)}_{1}$ (i.e., weighted according to cluster size). In the non-ergodic case, variables $\widetilde{X}^{\left(q\right)}$ and $\widetilde{\tau}^{\left(q\right)}$ refer to the global specific heat and the time scale measured on $\mathbf{L}^{\left(q\right)}_{1}$ under the infinite deceleration condition. Eqs. (25-26) actually serve as a special case of Eqs. (27-28) where there exists only one cluster. Please see Appendix G for derivations of these equations. Moreover, although the scaling relations in Eqs. (25-28) are derived in a case with $k=1$, we can also explore these scaling relations on $\mathbf{G}^{\left(q\right)}_{k}$ for $k>1$ (i.e., we directly replace $\mathbf{G}^{\left(q\right)}_{1}$ with $\mathbf{G}^{\left(q\right)}_{k}$ during calculating Eqs. (25-28)). In summary, we can verify if a system obeys scaling relations and further explore whether the SRG preserves scaling relations.
Eqs. (25-28) are derived for $q$-order interactions during renormalization and may divide from the original scaling relation of $p$-order interactions. These $p$-order interactions, initially corresponding to $\widetilde{X}^{\left(p\right)}$ and $\beta^{\left(p\right)}$, are renormalized according to $\widetilde{X}^{\left(q\right)}$ and $\beta^{\left(q\right)}$ in the SRG. In the case with $p<q$, we can measure the effects of $q$-order interactions on the scaling relation of $p$-order ones as
@@ V.1 Verification of the multi-order scale-invariance
In Appendix H, we implement our verification according to the behaviours of the Laplacian eigenvalue spectrum (the behaviours of degree distributions are also presented as auxiliary information). Based on the existence of invariant power-law forms of Laplacian eigenvalue spectra under the SRG transformation, we can classify systems into scale-invariant, weakly scale-invariant, and scale-dependent types on different orders.
@@
Apart from the above approach, there exists a more direct way to realize the verification. Because a scale-invariant system on $q$-order should invariably follow the scaling relation, being born to satisfy Eq. (27) is one of the necessary conditions of being scale-invariant. Consequently, we can first measure the deviations from the scaling relation of concerned systems before renormalization. Given significantly large deviations, the system can not be scale-invariant no matter how it behaves during renormalization. In Fig. 5, the measurement is conducted on multiple types of systems in each $k$-th iteration ($k\in\{1,2,3\}$) of the SRG (see Appendix A for random network definitions and see Appendix I for experiment details). As shown in Figs. 5(a-e), although the mean observations of $\left(\widetilde{X}^{\left(q\right)},\widetilde{\tau}^{\left(q\right)}\Big{\langle}\frac{\beta^{\left(q\right)}\left(i\right)+1}{\widetilde{\tau}^{\left(q\right)}\left(i\right)}\Big{\rangle}_{i}+\nu\right)$ averaged across all replicas collapse onto the scaling relation on every order and in each $k$-th iteration, the departures from the scaling relation measured by the standard deviations of $\widetilde{X}^{\left(q\right)}$ and $\widetilde{\tau}^{\left(q\right)}\Big{\langle}\frac{\beta^{\left(q\right)}\left(i\right)+1}{\widetilde{\tau}^{\left(q\right)}\left(i\right)}\Big{\rangle}_{i}+\nu$ in some systems can be non-negligible. Specifically, as suggested in Fig. 5(f), the systems whose pairwise interactions follow known scale-invariant structures, such as the triangular lattice or the random tree, do ensure that all their replicas follow the scaling relation with vanishing deviations in each $k$-th iteration. The replicas of the systems whose pairwise interactions follow the Barabási-Albert network, a structure with weak scale-invariant property (i.e., the Barabási-Albert network can be renormalized within a specific set of scales but is not strictly scale-invariant [10]), have relatively small deviations on the first three orders as well. The replicas of other kinds of systems whose pairwise interactions follow the Watts-Strogatz or the Erdos-Renyi networks [81] exhibit larger deviations from the scaling relation on most orders, suggesting the absence of mutli-order scale-invariance.
@@
In sum, the method reported in Fig. 5 can offer a convenient detection of scale-dependent systems in practice because scale-invariant systems are expect to satisfy Eq. (27) before and during renormalization. To seek a more precise verification, one should apply the approach in Appendix H.

### v7

@@ IV.2 Critical time scale
In the ergodic case, the setting of $\tau^{\left(q\right)}$ in the SRG (as we have mentioned in step (1)) is finished in $1$-st iteration and required to make the specific heat, an indicator of the transitions of diffusion processes, maximize (see similar ideas in Ref. [10]). In Appendix D, we explain why it is valid to define renormalization flows based on the specific heat. Specifically, we consider the $q$-order spectral entropy in the $1$-st iteration of the SRG (see Ref. [42] for the definition of spectral entropy)
@@
where each $\lambda^{\left(q\right)}_{1}$ is an eigenvalue of $\mathbf{L}^{\left(q\right)}_{1}$ and $\rho^{\left(q\right)}_{1}$ is the density operator derived from $\mathbf{L}^{\left(q\right)}_{1}$ following Eq. (11). Note that we have $\big{\langle}\lambda^{\left(q\right)}_{1}\big{\rangle}_{\rho}=\operatorname{tr}\left(\mathbf{L}^{\left(q\right)}_{1}\rho^{\left(q\right)}_{1}\right)$. The $q$-order specific heat is defined according to the first derivative of the $q$-order spectral entropy
@@
Please see Appendix E for detailed derivations. The value of $\tau^{\left(q\right)}$ is selected to ensure $\frac{\mathsf{d}}{\mathsf{d}\tau^{\left(q\right)}}X^{\left(q\right)}_{1}\left(\tau^{\left(q\right)}\right)=0$, which indicates an infinite deceleration condition of diffusion processes. See Fig. 3(c) for illustrations. Apart from the infinite deceleration condition, determining the time scale value $\widehat{\tau}^{\left(q\right)}$ used in the step (2) of the SRG also requires us to consider the high-order properties of diffusion. For convenience, we denote $\widetilde{\tau}^{\left(q\right)}$ as the value of $\tau^{\left(q\right)}$ under the infinite deceleration condition (i.e., there is $\frac{\mathsf{d}}{\mathsf{d}\tau^{\left(q\right)}}X^{\left(q\right)}_{1}\left(\tau^{\left(q\right)}\right)\big{|}_{\widetilde{\tau}^{\left(q\right)}}=0$). The optimal time scale for the SRG is set as
$\widehat{\tau}^{\left(q\right)}=\frac{2\varepsilon}{q+1}\widetilde{\tau}^{\left(q\right)},$ (26)
@@
which is the average subdivided time scale for each $1$-simplex in a $n$-simplex (i.e., the average number of adjacent $1$-simplices between arbitrary units $i$ and $j$ is $\frac{n+1}{2}$). Note that $\widehat{\tau}^{\left(q\right)}=\widetilde{\tau}^{\left(q\right)}$ holds for pairwise interactions. The value of $\varepsilon$ is used to correct the errors of time scale calculation when system units share ultra-dense interactions, which arise from the numerical bias of diffusion rate estimation in ultra-dense systems. The definition of $\varepsilon$ can be seen in Appendix F.
@@
The above procedure can be directly generalized to the non-ergodic case, where we treat each cluster $C_{i}$ as a network and derive a $\widehat{\tau}^{\left(q\right)}$ for it if it has not been assigned with a time scale yet. The time scale of a cluster does not change after it is assigned. See Fig. 3(c) for illustrations.
IV.3 Renormalization flow
After determining the time scale, the SRG proposed in Sec. IV.1 can be applied to renormalize interacting systems in a non-parametric manner. Please see Appendix M for the programmatic implementation of the SRG.
In Fig. 4(a), we explain how the SRG iterates. For an input system, we define $p$ and $q$ to realize the renormalization guided by different orders of interactions. Given each pair of $\left(p,q\right)$, we first search $p$- and $q$-simplices to represent the system on the $p$- and $q$-order (i.e., generate the high-order network sketches, $\mathbf{G}^{\left(p\right)}_{1}$ and $\mathbf{G}^{\left(q\right)}_{1}$, as shown in Fig. 1). Meanwhile, we derive the critical time scale, $\widehat{\tau}^{\left(q\right)}$, on $\mathbf{G}^{\left(q\right)}_{1}$ following Sec. IV.2. After the critical time scale is calculated, we iteratively realize the steps (1-6) introduced in Sec. IV.1 to renormalize the system once, twice, and so on. For instance, in the first iteration of the SRG, we follow steps (2-3) to find the Kadanoff blocks on the $q$-order (i.e., clusters in $\mathbf{H}^{\left(q\right)}_{1}$) and coarse grain the units in each Kadanoff block into a macro-unit on the $p$ and $q$-order to derive $\mathbf{G}^{\left(p\right)}_{2}$ and $\mathbf{G}^{\left(q\right)}_{2}$. Then, we follow steps (4-6) to calculate the associated high-order Laplacian operators $\mathbf{L}^{\left(p\right)}_{2}$ and $\mathbf{L}^{\left(q\right)}_{2}$. The obtained $\mathbf{G}^{\left(p\right)}_{2}$, $\mathbf{G}^{\left(q\right)}_{2}$, $\mathbf{L}^{\left(p\right)}_{2}$, and $\mathbf{L}^{\left(q\right)}_{2}$ serve as inputs for the SRG to run the second iteration to generate $\mathbf{G}^{\left(p\right)}_{3}$, $\mathbf{G}^{\left(q\right)}_{3}$, $\mathbf{L}^{\left(p\right)}_{3}$, and $\mathbf{L}^{\left(q\right)}_{3}$. In this manner, the iteration continues and a renormalization flow is gradually generated.
Figure 5: Large-scale instances of the renormalization flows of the simplex renormalization group (SRG). The first three iterations of renormalization on the systems whose pairwise interactions follow the random tree (a), the square lattice (b), the square lattice with a periodic boundary condition (c), the triangular lattice (d), the triangular lattice with a periodic boundary condition (e), the Watts-Strogatz networks (f-h), and the Erdos-Renyi networks (i-j) are presented, where we set $\left(p,q\right)=\left(1,1\right)$, $\left(p,q\right)=\left(1,2\right)$ and $\left(p,q\right)=\left(1,3\right)$ for (a-c, f, i), (d-e, g, j) and (h), respectively. Please see Appendix A for the definitions of these random network models. For convenience, the periodic boundary condition is abbreviated to the PBC. The absence of the PBC implies that the definitions of different lattices (i.e., square and triangular) only hold within system boundaries.
In Figs. 4(b-d), we mainly illustrate how the real space coarse graining in the SRG happens on different orders. We define $p=1$ and $q\in\{1,2,3\}$ to generate renormalization flows and visualize the changes caused by coarse graining the units within each Kadanoff block. Two important messages are conveyed by Fig. 4. First, whenever a Kadanoff block of $m$ units is find in $\mathbf{H}^{\left(q\right)}_{k}$, the size of $\mathbf{G}^{\left(p\right)}_{k+1}$ is reduced by $m-1$ (this property also holds for $\mathbf{G}^{\left(q\right)}_{k+1}$). If there is no Kadanoff block in $\mathbf{H}^{\left(q\right)}_{k}$ (i.e., no unit to coarse grain), the SRG arrives at its fixed point and the system remains unchanged during subsequent renormalization. Second, as shown in the comparison across different values of $q$, the system reduction during renormalization is always slighter given a larger value of $q$, which mainly arises from the sparser distribution of high-order interactions.
To demonstrate the meaning of distinguishing between $p$ and $q$ during renormalization, we compare between the cases with $p\neq q$ (i.e., renormalize the system on an order according to its properties on another order) and $p=q$ (i.e., equivalent to the situation where we do not distinguish between $p$ and $q$) in Appendix G. In Fig. 11, we observe significant differences in both real space (i.e., $\mathbf{G}^{\left(p\right)}_{k}$) and moment space (i.e., $\mathbf{L}^{\left(p\right)}_{k}$) between the cases of $p=q$ and $p\neq q$, demonstrating the effectiveness of supporting $p\neq q$ during renormalization. Therefore, by distinguishing between $p$ and $q$, the SRG is enabled to analyze more phenomena than classic frameworks with $p\equiv q$. Meanwhile, these results qualitatively suggest the effects of higher-order interactions on lower-order interactions during renormalization (i.e., in the case with $p\neq q$), whose quantification is explored in our subsequent analysis.
In Fig. 5, we present more instances of the renormalization flow generated by the SRG. We apply the SRG on synthetic interacting systems whose pairwise interactions follow the random tree (RT), the square lattice (SL), the triangular lattice (TL), the Watts-Strogatz networks (WS), and the Erdos-Renyi networks (ER). Among these random networks, the square and triangular lattices can be defined with or without the periodic boundary condition. Please see Appendix A for the definitions of these random networks. As shown in Fig. 5, the key structural properties of RT, SL, and TL are principally preserved during renormalization (i.e., we can see qualitative similarities in $\mathbf{G}^{\left(p\right)}_{k}$ across different iterations) while the SRG implemented on WS and ER gradually reveals specific latent structures that are dissimilar to the original ones. Among these instances, the lattice systems with periodic boundary conditions are observed to make the renormalization ineffective. This is a phenomenon caused by the identical properties of all units under the periodic boundary condition (i.e., all units are exactly the same as each other on any $q$-order). Because the renormalizaton procedure in the SRG begins with distinguishing between short- and long-range interactions, the SRG can effectively reduce a system only when units and interactions are not identical. Otherwise, it is impossible to tell which interactions are long-range and which are short-range. In Fig. 5(c) and Fig. 5(e), the SRG does not coarse grain systems because it discover no short-range interaction. This phenomenon does not cause conflicts in our work because a lattice system with identical units and interactions can be treated as an extreme case of self-similarity to certain extent (i.e., each sub-system is identical to the whole system) and can be invariant for coarse graining. To avoid this trivial phenomenon in our analysis, all the lattice systems implemented in our subsequent experiments (e.g., Figs. 6-7) are defined without the periodic boundary condition.
IV.4 Scaling relation and high-order effects
@@
where $\big{\langle}\lambda^{\left(q\right)}_{1}\big{\rangle}_{\widetilde{\rho}}=\operatorname{tr}\left(\mathbf{L}^{\left(q\right)}_{1}\widetilde{\rho}^{\left(q\right)}\right)$ and the constant $\mu$ is obtained when we solve Eq. (25) as a differential equation. If the Laplacian eigenvalue spectrum (i.e., the density of the eigenvalues of $\mathbf{L}^{\left(q\right)}_{1}$) follows a general power-law form $\operatorname{Prob}\left(\lambda^{\left(q\right)}_{1}\right)\sim\left(\lambda^{\left(q\right)}_{1}\right)^{\beta^{\left(q\right)}}$, we can obtain a scaling relation
@@
whose derivations are presented in Appendix H. The parameter $\nu$ is a specific constant. For a system with scale-invariance property on the $q$-order, the power-law form of the Laplacian eigenvalue spectrum holds and should be invariant under the transformation of the SRG. Therefore, we expect to see an approximately constant $\widetilde{X}^{\left(q\right)}$ in a certain interval of time scale (i.e., the infinite deceleration condition), accompanied by a fixed $\beta^{\left(q\right)}$ during renormalization. For a system without scale-invariance, the scaling relation does not hold true.
Figure 6: Scaling relation of the simplex renormalization group. (a-e) We verify the scaling relation on the systems whose pairwise interactions follow a Barabási-Albert network (BA, $c=3$), a Watts-Strogatz network (WS, each unit initially has $8$ neighbors and edges are rewired according to a probability of $0.3$), and an Erdos-Renyi network (ER, two units share an edge with a probability of $0.08$), the triangular lattice (TL, with the interactions of the first two orders), and the random tree (RT, with only pairwise interactions), respectively. Scatters denote the mean values of $\left(\widetilde{X}^{\left(q\right)},\widetilde{\tau}^{\left(q\right)}\Big{\langle}\frac{\beta^{\left(q\right)}\left(i\right)+1}{\widetilde{\tau}^{\left(q\right)}\left(i\right)}\Big{\rangle}_{i}+\nu\right)$ averaged across all replicas on each order and error bars denote standard deviations. See the definitions of random networks in Appendix A. (f) The standard deviations compared across different systems. (g-i) The high-order effects measured on the replicas in (a-c), where each color area denotes the distribution of high-order effects in the $k$-iteration of the SRG ($k\in\{1,2,3\}$).
To generalize our analysis to the non-ergodic case, we respectively derive $\widetilde{\tau}^{\left(q\right)}\left(i\right)$ and $\beta^{\left(q\right)}\left(i\right)$ for each cluster $C_{i}$ of $\mathbf{G}^{\left(q\right)}_{1}$ following Eqs. (27-28). These definitions enable us to obtain
@@
where $\langle\cdot\rangle_{i}$ denotes weighted averaging across all clusters of $\mathbf{G}^{\left(q\right)}_{1}$ (i.e., weighted according to cluster size). In the non-ergodic case, variables $\widetilde{X}^{\left(q\right)}$ and $\widetilde{\tau}^{\left(q\right)}$ refer to the global specific heat and the time scale measured on $\mathbf{L}^{\left(q\right)}_{1}$ under the infinite deceleration condition. Eqs. (27-28) actually serve as a special case of Eqs. (29-32) where there exists only one cluster. Please see Appendix I for derivations of these equations. Moreover, although the scaling relations in Eqs. (27-32) are derived in a case with $k=1$, we can also explore these scaling relations on $\mathbf{G}^{\left(q\right)}_{k}$ for $k>1$ (i.e., we directly replace $\mathbf{G}^{\left(q\right)}_{1}$ with $\mathbf{G}^{\left(q\right)}_{k}$ during calculating Eqs. (27-32)). In summary, we can verify if a system obeys scaling relations and further explore whether the SRG preserves scaling relations.
Eqs. (27-32) are derived for $q$-order interactions during renormalization and may divide from the original scaling relation of $p$-order interactions. These $p$-order interactions, initially corresponding to $\widetilde{X}^{\left(p\right)}$ and $\beta^{\left(p\right)}$, are renormalized according to $\widetilde{X}^{\left(q\right)}$ and $\beta^{\left(q\right)}$ in the SRG. In the case with $p<q$, we can measure the effects of $q$-order interactions on the scaling relation of $p$-order ones as
@@ V.1 Verification of the multi-order scale-invariance
In Appendix J, we implement our verification according to the behaviours of the Laplacian eigenvalue spectrum (the behaviours of degree distributions are also presented as auxiliary information). Based on the existence of invariant power-law forms of Laplacian eigenvalue spectra under the SRG transformation, we can classify systems into scale-invariant, weakly scale-invariant, and scale-dependent types on different orders.
@@
Apart from the above approach, there exists a more direct way to realize the verification. Because a scale-invariant system on the $q$-order should invariably follow the scaling relation, being bound to satisfy Eq. (29) is one of the necessary conditions of being scale-invariant. Consequently, we can first measure the deviations from the scaling relation of concerned systems before renormalization. Given significantly large deviations, the system can not be scale-invariant no matter how it behaves during renormalization. In Fig. 6, the measurement is conducted on multiple types of systems in each $k$-th iteration ($k\in\{1,2,3\}$) of the SRG (see Appendix A for random network definitions and see Appendix K for experiment details). As shown in Figs. 6(a-e), although the mean observations of $\left(\widetilde{X}^{\left(q\right)},\widetilde{\tau}^{\left(q\right)}\Big{\langle}\frac{\beta^{\left(q\right)}\left(i\right)+1}{\widetilde{\tau}^{\left(q\right)}\left(i\right)}\Big{\rangle}_{i}+\nu\right)$ averaged across all replicas collapse onto the scaling relation on every order and in each $k$-th iteration, the departures from the scaling relation measured by the standard deviations of $\widetilde{X}^{\left(q\right)}$ and $\widetilde{\tau}^{\left(q\right)}\Big{\langle}\frac{\beta^{\left(q\right)}\left(i\right)+1}{\widetilde{\tau}^{\left(q\right)}\left(i\right)}\Big{\rangle}_{i}+\nu$ in some systems can be non-negligible. Specifically, as suggested in Fig. 6(f), the systems whose pairwise interactions follow known scale-invariant structures, such as the triangular lattice or the random tree, do ensure that all their replicas follow the scaling relation with vanishing deviations in each $k$-th iteration. The replicas of the systems whose pairwise interactions follow the Barabási-Albert network, a structure with weak scale-invariant property (i.e., the Barabási-Albert network can be renormalized within a specific set of scales but is not strictly scale-invariant [10]), have relatively small deviations on the first three orders as well. The replicas of other kinds of systems whose pairwise interactions follow the Watts-Strogatz or the Erdos-Renyi networks [80] exhibit larger deviations from the scaling relation on most orders, suggesting the absence of mutli-order scale-invariance.
@@
In sum, the method reported in Fig. 6 can offer a convenient detection of scale-dependent systems in practice because scale-invariant systems are expect to satisfy Eq. (29) before and during renormalization. To seek a more precise verification, one should apply the approach in Appendix J.